OpenStudy (nilankshi):

A bus decreases its speed from 80 km/h in 5s find acceleration of the bus

OpenStudy (anonymous):

80 km/h to 0km/h?

OpenStudy (anonymous):

because then its (Vf-Vi)/time=Acc. so (80-0)/5=Acc. Acc=16

OpenStudy (anonymous):

V(t)=U+at v(t)=0 U=80km/hr=22.22 m/s t=5s \[V(t)=U+at\] \[0=22.22+(a*5)\] \[a=-22.22/5\] \[a-4.44m/s^2\]

OpenStudy (anonymous):

i think this is wrong